OpenStudy (anonymous):

What are all solutions of 2x2 - 5x = x2 - 6x + 2 ? A) 1 B) 3 C) -2 D) -2,1 E) 1,3

OpenStudy (anonymous):

are any of the numbers supposed to be squared?

OpenStudy (johnweldon1993):

2x² - 5x = x² - 6x + 2 ? First thing you want to do is combine like terms....so try subtract x² from both sides of the equation...and try adding 6x to both sides as well... this will leave you with \[\large x^2 + x = 2\] Can you do this now? you can either complete the square or use the quadratic formula... to solve this...

OpenStudy (anonymous):

no

OpenStudy (johnweldon1993):

Lets use the quadratic formula.....we need to set our equation = to 0 and solve for 'x' \[\large x^2 + x - 2 = 0\] you have a form of \[\large ax^2 + bx + c\] so your a = 1 b = 1 c = -2 use the equation \[\large \frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\] plug those into the equation and solve

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

i got 1,-2